If x=-3 is a solution of the equation x^2=6x+do where d is a constant what is another value of x that satisfies the equation?
(Plz show step by step solution)
-5
-9
9
6

Answer by ikleyn(52781)

.
x^2 = 6x + d    ====>


x^2 - 6x - d = 0.


According to Vieta's theorem, the sum of the roots to this equation is equal to the coefficient at x with the opposite sign:


alpha + beta = 6  ====>


beta = 6 - alpha = 6 - (-3) = 9.


Answer.  The other root is 9.

Solved.

==============

On Vieta's theorem, see the lesson
    - Using Vieta's theorem to solve qudratic equations and related problems
in this site.